\begin{thebibliography}{10}

\bibitem{Ahn1981}
Byong~Hun Ahn.
\newblock Solution of nonsymmetric linear complementarity problems by iterative
  methods.
\newblock {\em Journal of optimization Theory and Applications}, 33:175--185,
  1981.

\bibitem{Bai2010}
Zhong-Zhi Bai.
\newblock Modulus-based matrix splitting iteration methods for linear
  complementarity problems.
\newblock {\em Numerical Linear Algebra with Applications}, 17(6):917--933,
  2010.

\bibitem{Ben1975}
Alain Bensoussan and Jacques-Louis Lions.
\newblock Nouvelles m{\'e}thodes en contr{\^o}le impulsionnel.
\newblock {\em Applied Mathematics and Optimization}, 1:289--312, 1975.

\bibitem{Cottle1992}
Richard~W Cottle, Jong-Shi Pang, and Richard~E Stone.
\newblock The linear complementarity problem. academic pr ess.
\newblock {\em Inc., Boston, MA}, 5, 1992.

\bibitem{Ferris2011}
Michael~C Ferris, Olvi~L Mangasarian, and Jong-Shi Pang.
\newblock {\em Complementarity: Applications, algorithms and extensions},
  volume~50.
\newblock Springer Science \& Business Media, 2001.

\bibitem{Hong2016}
Jun-Tao Hong and Chen-Liang Li.
\newblock Modulus-based matrix splitting iteration methods for a class of
  implicit complementarity problems.
\newblock {\em Numerical Linear Algebra with Applications}, 23(4):629--641,
  2016.

\bibitem{Kumarnote}
Bharat Kumar, Arup~K Das, et~al.
\newblock A note on fixed point method and linear complementarity problem.
\newblock {\em Journal of Numerical Analysis and Approximation Theory}, 2023.

\bibitem{Kumaron}
Bharat Kumar, Arup~Kumar Das, et~al.
\newblock On general fixed point method based on matrix splitting for solving
  linear complementarity problem.
\newblock {\em Journal of Numerical Analysis and Approximation Theory},
  51(2):189--200, 2022.

\bibitem{Kumarpro}
Bharat Kumar, Deepmala, and AK~Das.
\newblock Projected fixed point iterative method for large and sparse
  horizontal linear complementarity problem.
\newblock {\em Indian Journal of Pure and Applied Mathematics}, pages 1--10,
  2023.

\bibitem{Kumarmore}
Bharat Kumar, Deepmala, A~Dutta, and AK~Das.
\newblock More on matrix splitting modulus-based iterative methods for solving
  linear complementarity problem.
\newblock {\em OPSEARCH}, pages 1--18, 2023.

\bibitem{Lemke1964}
Carlton~E Lemke and Joseph~T Howson, Jr.
\newblock Equilibrium points of bimatrix games.
\newblock {\em Journal of the Society for industrial and Applied Mathematics},
  12(2):413--423, 1964.

\bibitem{Mangasarian1977}
OL~Mangasarian.
\newblock Solution of symmetric linear complementarity problems by iterative
  methods.
\newblock {\em Journal of Optimization Theory and Applications},
  22(4):465--485, 1977.

\bibitem{Mezzadri2020}
Francesco Mezzadri and Emanuele Galligani.
\newblock Modulus-based matrix splitting methods for horizontal linear
  complementarity problems.
\newblock {\em Numerical Algorithms}, 83(1):201--219, 2020.

\bibitem{Murty1988}
Katta~G Murty and Feng-Tien Yu.
\newblock {\em Linear complementarity, linear and nonlinear programming},
  volume~3.
\newblock Citeseer, 1988.

\bibitem{Noor1988}
Muhammad~Aslam Noor.
\newblock Fixed point approach for complementarity problems.
\newblock {\em Journal of Mathematical Analysis and Applications},
  133(2):437--448, 1988.

\bibitem{Xie2016}
Shui-Lian Xie, Hong-Ru Xu, and Jin-Ping Zeng.
\newblock Two-step modulus-based matrix splitting iteration method for a class
  of nonlinear complementarity problems.
\newblock {\em Linear Algebra and Its Applications}, 494:1--10, 2016.

\end{thebibliography}

\documentclass[12pt]{article}
\usepackage{enumerate}
\usepackage{amsmath}
\usepackage{amsthm}
\usepackage{amssymb}
\usepackage{algorithm}
\usepackage[bookmarksnumbered,  plainpages]{hyperref}

\newtheorem{thm}{Theorem}[section]

\newtheorem{lem}{Lemma}[section]

\newtheorem{prop}[thm]{Proposition}
\theoremstyle{definition}

\theoremstyle{remark}
\newtheorem{rem}{Remark}[section]
\begin{document}
	
	\title{{More on Equivalent Formulation of  Implicit Complementarity  Problem}}
	
	\author{Bharat kumar$^{a,1}$, Deepmala$^{a,2}$ and A.K. Das $^{b,3}$\\
		\emph{\small PDPM-Indian Institute of Information Technology, Design and Manufacturing,}\\
		\emph{\small Jabalpur - 482005 (MP), India}\\
		\emph{\small $^{b}$Indian Statistical Institute, 203 B.T. Road, Kolkata - 700108, India }\\
		\emph{\small $^1$Email:bharatnishad.kanpu@gmail.com , $^2$Email: dmrai23@gmail.com}\\		\emph{\small $^3$Email: akdas@isical.ac.in}}
	\date{}
	\maketitle
	%%==================================%%
	%% sample for unstructured abstract %%
	%%==================================%%
	
	\abstract{\noindent  This article presents an equivalent formulation  of the implicit complementarity problem. We  demonstrate that solution of the equivalent  formulation is equivalent to the solution of the implicit complementarity problem. Moreover, we provide another equivalent formulation of the implicit complementarity problem using a strictly increasing function.
	 }

	\noindent \textbf{Keywords.} Implicit complementarity problem, equivalent form, strictly increasing function.\\
	
	\noindent \textbf{Mathematics Subject Classification.}  90C33, 90C30.\\
	%%\pacs[JEL Classification]{D8, H51}
	
	%%\pacs[MSC Classification]{35A01, 65L10, 65L12, 65L20, 65L70}
	
	\maketitle
	
	\section{Introduction}\label{sec1}
Bensoussan et al. \cite{Ben1975} presented the implicit complementarity problem (ICP).  	 ICPs are a class of mathematical optimization problems that solve a system of nonlinear equations that includes both complementary conditions and equality or inequality constraints. ICPs can be formulated as follows: \\Consider the matrix  $\mathcal{A} \in \mathbb{R}^{n\times n}$  and  the vector  $b \in \mathbb{R}^n$,  the implicit complementarity problem denoted as ICP$( \mathcal{A}, b, f )$ is to find the solution $ r \in \mathbb{R}^n $ to the following system:
\begin{equation}\label{eq0}
	\mathcal{H}(r)=r-f(r)\geq 0,~~~  \mathcal{F}(r)=\mathcal{A}r+b \geq 0,~~~ \mathcal{H}(r)^T\mathcal{F}(r)=0,	
\end{equation}
where $f(r)$ is a mapping from $\mathbb{R}^n$ to $\mathbb{R}^n$.
ICPs arise in many different areas of mathematics and science, including economics, engineering, physics and computer science. For details see,  \cite{Lemke1964},  \cite{Ferris2011} and \cite{Murty1988}.\\

\noindent One of the most popular techniques for developing fast and affordable iterative algorithms is the equivalent formulation of the LCP$(\mathcal{A}, b)$ as an equation that's solution must be the same.  Mangasarian offered an equivalent forms    of  LCP$(\mathcal{A}, b)$  in \cite{Mangasarian1977} and described as 
\begin{equation*}
	r = (r- \omega \Omega(\mathcal{A}r + b))_+,
\end{equation*}
where  $r_+ \in \mathbb{R}^n$, $(r_+)_i = max\{0, r_i\}$ and  $\Omega \in \mathbb{R}^{n\times n}$ is a positive diagonal matrix.   The LCP $(\mathcal {A}, b)$ is described in an analogous form and several iteration techniques are given by Bai in \cite{Bai2010}. For more details on equivalent form of  LCPs and related iteration methods see, \cite{Ahn1981}, \cite{Hong2016}, \cite{Kumaron}, \cite{Kumarmore}, \cite{Noor1988}, \cite{Kumarpro}, \cite{Kumarnote} and  \cite{Cottle1992}. The concept of equivalent formulation has also been used effectively for other complementarity problems, like  implicit complementarity problem \cite{Xie2016} and  \cite{Hong2016} and  horizontal linear complementarity problem \cite{Mezzadri2020}.
Motivated by the works of Mangasarian \cite{Mangasarian1977}, we present an equivalent form of ICP. 
\noindent  The article is presented as follows:   we present an equivalent form of ICP  and provide some  conditions for the solution of ICP in Section 2. The  conclusion is given in Section 3.

	%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%
	\section{Main results}\label{sec2}
\noindent In the following, we provide an equivalent expression of the implicit complementarity problem.
\begin{thm}\label{thm1}
	Suppose  $\mathcal{A} \in \mathbb{R}^{n \times n}$   and $b \in \mathbb{R}^{n}$. Then $r^* \in \mathbb{R}^{n \times n }$  be the solution of   ICP$( \mathcal{A}, b, f)$ $\iff$ $\mathcal{R}(r^*)=0$, where  $\mathcal{R} : \mathbb{R} \rightarrow \mathbb{R}$ is defined as 
	\begin{equation}\label{E1}
		\mathcal{R}(r)=	\mathcal{H}(r)-(\mathcal{H}(r)-(\mathcal{A}r+b))_{+}.		
	\end{equation}
\end{thm}
\begin{proof} Suppose $\mathcal{R}(r^*)=0$, it follows that $(\mathcal{H}(r^*)-(\mathcal{H}(r^*)-(\mathcal{A}r^*+b))_{+})=0$. This implies that 
	\begin{equation}\label{eq1}
		\mathcal{H}(r^*)=(\mathcal{H}(r^*)-(\mathcal{A}r^*+b))_{+}.
	\end{equation}
	Component-wise, we  consider two cases:\\
	\textbf{Case 1.} when $\mathcal{H}_{i}(r^*) \geq (\mathcal{A}r^*+b)_{i}$, where $\mathcal{H}_{i}(r^*)$ denotes the $i^{th}$ component of the $\mathcal{H}(r^*)$. Then   Eq.  (\ref{eq1}) can be written as 
	\begin{equation*}
		\mathcal{H}_{i}(r^*)=\mathcal{H}_{i}(r^*)-(\mathcal{A}r^*+b)_{i}.
	\end{equation*}
	It follows that, $$\mathcal{F}_{i}(r^*)=(\mathcal{A}r^*+b)_{i}=0.$$\\
	\textbf{Case 2.}  when $\mathcal{H}_{i}(r^*) ~\textless ~(\mathcal{A}r^*+b)_{i}$ $\implies$  $((\mathcal{A}r^*+b)_{i}- \mathcal{H}_{i}(r^*))\ \textless ~0$. Then, we get  
	\begin{equation*}
		\mathcal{H}_{i}(r^*)=0.
	\end{equation*}
	From case (1) and case (2),  $\mathcal{H}_{i}(r^*)\mathcal{F}_{i}(r^*)=0 ~\forall ~ i $  $\implies$ $\mathcal{H}(r^*)^{T}\mathcal{F}(r^*)=0 $\\
	Conversely, let $r^*$ be the solution of  system (\ref{eq0}). By  complementary condition  of ICP,  either $\mathcal{H}_{i}(r^*)=0$ or $\mathcal{F}_{i}(r^*)=0 ~\forall ~i$.\\	
	Component-wise, we  consider two cases:\\
	\textbf{Case 1.} If $\mathcal{H}_{i}(r^*)=0$  and $\mathcal{F}_{i}(r^*) \textgreater 0$,  Eq. (\ref{E1}) becomes 
	$\mathcal{R}_{i}(r^*)=(-(-(\mathcal{A}r^*+b))_{+})$
	$\implies $ $\mathcal{R}_{i}(r^*)=0.$\\
	\textbf{Case 2.} If $\mathcal{H}_{i}(r^*) \textgreater 0$  and $\mathcal{F}_{i}(r^*) = 0$, then 
	$\mathcal{R}_{i}(r^*)=(\mathcal{H}_{i}(r^*)-(\mathcal{H}_{i}(r^*))_{+})$
	$\implies$ $\mathcal{R}_{i}(r^*)=0.$\\
	From case (1) and case (2), we get  $\mathcal{R}_{i}(r^*)=0 ~ \forall ~ i.$ Then  $\mathcal{R}(r^*)=0$
\end{proof}
\begin{rem} Let $\mathcal{S}(r)=\mathcal{H}(r)- \mathcal{R}(r)$, then 
	$r^*$ is the solution of  ICP$( \mathcal{A}, b, f )$ $\iff$  $\mathcal{S}(r^*)= \mathcal{H}(r^*)$.	
\end{rem}
\begin{lem}
	Suppose   $\mathcal{A} \in \mathbb{R}^{n \times n}$  and $b \in \mathbb{R}^{n}$. Let  $\Omega_{1}, \Omega_{2} \in \mathbb{R}^{n \times n}$  be  two positive diagonal matrices and  define the mapping $\bar{\mathcal{R}}(r)= (\Omega_{1}\mathcal{H}(r)- (\Omega_{1}\mathcal{H}(r)-\Omega_{2}(\mathcal{A}r+b))_{+})$. Then  $r^* $ is the solution of ICP$( \mathcal{A}, b, f )$ $\iff$   $\bar{\mathcal{R}}(r^*)=0$ 
\end{lem}
\begin{proof} Suppose $\bar{\mathcal{R}}(r^*)=0$, it follows that $$(\Omega_{1}\mathcal{H}(r^*)-(\Omega_{1}\mathcal{H}(r^*)-\Omega_{2}(\mathcal{A}r^*+b))_{+})=0.$$ Then  we write  
	\begin{equation}\label{eq2}
		\Omega_{1}\mathcal{H}(r^*)=(\Omega_{1}\mathcal{H}(r^*)-\Omega_{2}(\mathcal{A}r^*+b))_{+}.
	\end{equation}
	Component-wise, we  consider two cases:\\
	\textbf{Case 1.} when $(\Omega_{1}\mathcal{H}(r^*))_{i} \geq (\Omega_{2}(\mathcal{A}r^*+b))_{i}$, then   Eq.  (\ref{eq1}) becomes 
	\begin{equation*}
		(\Omega_{1}\mathcal{H}(r^*))_{i}=(\Omega_{1}\mathcal{H}(r^*))_{i}-(\Omega_{2}(\mathcal{A}r^*+b))_{i}.
	\end{equation*}
	Then,  
	$(\Omega_{2}(\mathcal{A}r^*+b))_{i}=0$ $\implies$
	$\mathcal{F}_{i}(r^*)=(\mathcal{A}r^*+b)_{i}=0$.\\
	\textbf{Case 2.}  when $(\Omega_{1}\mathcal{H}(r^*))_{i} ~\textless ~ (\Omega_{2}(\mathcal{A}r^*+b))_{i}$, this implies that $((\Omega_{2}(\mathcal{A}r^*+b))_{i}- (\Omega_{1}\mathcal{H}(r^*))_{i})\ \textless 0$. Then, we get  
	\begin{equation*}
		(\Omega_{1}\mathcal{H}(r^*))_{i}=0 \implies \mathcal{H}_{i}(r^*)=0,
	\end{equation*}
	from case (1) and case (2), we obtain  $\mathcal{H}_{i}(r^*)\mathcal{F}_{i}(r^*)=0 ~\forall ~i.  $ Hence,  $\mathcal{H}(r^*)^{T}\mathcal{F}(r^*)=0.  $\\
	Conversely, let $r^*$ be the solution of  Eq.  (\ref{eq0}). Then \\
	component-wise, we  consider two cases:\\
	\textbf{Case 1.} when $\mathcal{H}_{i}(r^*)=0$  and $\mathcal{F}_{i}(r^*) ~\textgreater~ 0$, then
	$\bar{\mathcal{R}}_{i}(r^*)=-(-(\Omega_{2}(\mathcal{A}r^*+b))_{i})_{+}$
	$\implies$ $\bar{\mathcal{R}}_{i}(r^*)=0.$\\
	\textbf{Case 2.} when $\mathcal{H}_{i}(r^*) ~\textgreater~ 0$  and $\mathcal{F}_{i}(r^*) = 0$, then 
	$\bar{\mathcal{R}}_{i}(r^*)=((\Omega_{1}\mathcal{H}(r^*))_{i}-((\Omega_{1}\mathcal{H}(r^*))_{i})_{+}).$\\
	Thus, $\bar{\mathcal{R}}_{i}(r^*)=0 ~\forall ~i.$
	From case (1) and case (2), we obtain  $\bar{\mathcal{R}}(r^*)=0.$
\end{proof}
\begin{prop}
	Suppose  $\mathcal{A} \in \mathbb{R}^{n \times n}$  and $b \in \mathbb{R}^{n}$. Let  $\delta : \mathbb{R} \rightarrow \mathbb{R}$ be any strictly increasing function  such that $\delta(0)=0$. Then   $r^* $ is the solution of  ICP$( \mathcal{A}, b, f )$ $\iff$ $\mathcal{G}(r^*)=0$,  $\mathcal{G}$ is the function from  $\mathbb{R}$ to  $\mathbb{R}$, given as 
	\begin{equation}\label{eq3}
		\mathcal{G}_{i}(r)=\delta(\lvert (\mathcal{A}r+ b )_{i}-\mathcal{H}_{i}(x)\rvert)-\delta((\mathcal{A}r+b)_{i})-\delta(\mathcal{H}_{i}(x)), ~~ i=1,2,\ldots,n.		
	\end{equation}
\end{prop}
\begin{proof} For some $i$, let $\mathcal{H}_{i}(r^*)~\textless~ 0$.  Then it follows that
	\begin{equation}\label{eq4}
		0 ~\textgreater ~\delta(\mathcal{H}_{i}(r^*))= \delta (\mathcal{F}_{i}(r^*)-\mathcal{H}_{i}(r^*))-\delta(\mathcal{F}_{i}(r^*)) \geq -\delta(\mathcal{F}_{i}(r^*)).
	\end{equation} 	
	Thus, $\mathcal{F}_{i}(r^*) \textgreater 0 $ and  $\mathcal{F}_{i}(r^*)-\mathcal{H}_{i}(r^*) ~\textgreater~ \mathcal{F}_{i}(r^*) ~ \textgreater~ 0$. This implies that  
	\begin{equation}\label{eq5}
		\delta (|\mathcal{F}_{i}(r^*)-\mathcal{H}_{i}(r)|)=  \delta (\mathcal{F}_{i}(r^*)-\mathcal{H}_{i}(r^*)) \textgreater \delta (\mathcal{F}_{i}(r)). \end {equation}
		From inequalities  (\ref{eq4}) and   (\ref{eq5}), we get
		$\mathcal{G}(r^*)~\textgreater~ 0$, this is the contradiction.\\
		If $\mathcal{H}_{i}(r^*)~\textgreater ~ 0$  and $\mathcal{F}_{i}(r^*) ~ \textgreater ~ 0$ for some $i$.
		Then, we consider two possibilities:\\
		1. when $\mathcal{H}_{i}(r^*) \textgreater \mathcal{F}_{i}(r^*)$, then 
		\begin{equation*}
			\delta (|\mathcal{F}_{i}(r^*)-\mathcal{H}_{i}(r^*)|)=  \delta (\mathcal{H}_{i}(r^*)-\mathcal{F}_{i}(r^*)) \textless \delta (\mathcal{H}_{i}(r^*)). \end {equation*}
			Then 
			$	\delta (|\mathcal{F}_{i}(r^*)-\mathcal{H}_{i}(r^*)|) - \delta (\mathcal{H}_{i}(r^*)) \textless  0$.
			This implies that 
			\begin{equation}\label{eq6}
				\begin{split}
					\mathcal{G}_{i}(r^*) \textless 0,		
				\end{split}
			\end{equation}
			2. when $\mathcal{F}_{i}(r^*) \textgreater \mathcal{H}_{i}(r^*)$, then 
			\begin{equation*}
				\delta (|\mathcal{F}_{i}(r^*)-\mathcal{H}_{i}(r^*)|)=  \delta (\mathcal{F}_{i}(r^*)-\mathcal{H}_{i}(r^*)) \textless \delta (\mathcal{F}_{i}(r^*)). \end {equation*}
				Then 
				$\delta (|\mathcal{F}_{i}(r^*)-\mathcal{H}_{i}(r^*)|) - \delta (\mathcal{F}_{i}(r^*)) \textless  0$.
				This implies that 
				\begin{equation}\label{eq7}
					\begin{split}
						\mathcal{G}_{i}(r^*) \textless 0.	
					\end{split}
				\end{equation}
				From inequalities (\ref{eq6}) and   (\ref{eq7}), we must have $\mathcal{G}_{i}(r^*) \textless 0$, again a contradiction.  Therefore, $r^*$ solve the   ICP$( \mathcal{A}, b, f)$.\\
				Conversely, let $r^* $ be the solution of   ICP$( \mathcal{A}, b, f)$, then  either $\mathcal{H}_{i}(r^*)=0$ or $\mathcal{F}_{i}(r^*)=0 ~\forall ~i$.\\
				Suppose $\mathcal{H}_{i}(r^*)=0$  and  $\mathcal{F}_{i}(r^*)\textgreater 0$, Then 
				\begin{equation*}
					\mathcal{G}_{i}(r^*)=\delta(\lvert \mathcal{F}_{i}(r^*) |)-\delta(\mathcal{F}_{i}(r^*)).
				\end{equation*} 
				This implies that $	\mathcal{G}_{i}(r^*)=0$.\\
				\noindent 	When $\mathcal{F}_{i}(r^*)=0$  and  $\mathcal{H}_{i}(r^*)\textgreater 0$, Then 
				\begin{equation*}
					\mathcal{G}_{i}(r^*)=\delta(\lvert  -\mathcal{H}_{i}(r^*)|)-\delta(\mathcal{H}_{i}(r^*)).
				\end{equation*} 
				This implies that $	\mathcal{G}_{i}(r^*)=0 ~ \forall~ i$. Therefore, $	\mathcal{G}(r^*)=0$
			\end{proof}	
\section{Conclusion} 
This article presents an equivalent formulation of the implicit complementarity problem. We introduce a equivalent  formulation of the implicit complementarity problem and demonstrate that solution of the equivalent  formulation is equal to the solution of the implicit complementarity problem. Moreover, By using strictly increasing function $\delta $, an  equivalent form of the implicit complementarity problem is provided. 
	
	\subsection*{Acknowledgments}
	The author, Bharat Kumar is thankful to the University Grants Commission (UGC), Government of India, under the SRF fellowship, Ref. No.: 1068/(CSIR-UGC NET DEC. 2017).
	%\section*{Authors' information}%% if any
	%Text for this section\ldots
	%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%
	%%                  The Bibliography                       %%
	%%                                                         %%
	%%  Bmc_mathpys.bst  will be used to                       %%
	%%  create a .BBL file for submission.                     %%
	%%  After submission of the .TEX file,                     %%
	%%  you will be prompted to submit your .BBL file.         %%
	%%                                                         %%
	%%                                                         %%
	%%  Note that the displayed Bibliography will not          %%
	%%  necessarily be rendered by Latex exactly as specified  %%
	%%  in the online Instructions for Authors.                %%
	%%                                                         %%
	%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%
	% if your bibliography is in bibtex format, use those commands:
	%\bibliographystyle{plain} % Style BST file (bmc-mathphys, vancouver, spbasic).
	%	\bibliography{bibfile}      % Bibliography file (usually '*.bib' )
	% for author-year bibliography (bmc-mathphys or spbasic)
	% a) write to bib file (bmc-mathphys only)
	% @settings{label, options="nameyear"}
	% b) uncomment next line
	%\nocite{label}
	% or include bibliography directly:
	\bibliographystyle{plain}
	\bibliography{p11}

\end{document}